Alex Adam Good A Good A Good B 0 160 30 Good B 300 225 150 75 0 120 80 25 50 75 100 60 90 40 120 a) Who has the comparative ab) Who has the comparative advantage in the production of good B?
a. Alex

b. Adam

c. Both Alex and Adam

d. Neither Alex nor Adam
Answer #1

Ans) Comparative advantage is when one can produce any good with less opportunity cost.

For Alex÷

Opportunity cost of good A = good B sacrificed ÷ good A gained = 75÷25 = 3

Opportunity cost of good B = good A sacrificed ÷ good B gained = 25÷75 = 0.33

[Difference between any point in producing good A is 25 and difference between any point in producing good B ks 75]

For Adam÷

Opportunity cost of producing good A = good B sacrificed ÷ good A gained = 40÷30 = 1.33

Opportunity cost of producing good B= good A sacrificed ÷ good B gained = 30÷40 = 0.75

[Difference between any point in producing good A is 30 and difference between any point in producing good B is 40]

Here, opportunity cost of producing good A is less for Adam, so he has comparative advantage in producing good A.

Here, opportunity cost of producing good B is less for Alex, so he has comparative advantage in producing good B.

1) Option b

2) Option a